OpenStudy (anonymous):

find the inverse laplace of F(s) = e^(-3s)/(s^2).

zepdrix (zepdrix):

So this relates to the unit step function. Ummm, I'm trying to remember the formula...\[\large \mathcal L^{-1}\left[e^{-as}F(s)\right]\quad =\quad f(t-a)\;u(t-a)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but it's u(t - a)f(t-a)

zepdrix (zepdrix):

How is that different from what I wrote? :O

zepdrix (zepdrix):

Oh based on the fact that I wrote the exponential first? :) fair enough!

OpenStudy (anonymous):

yeah haha

zepdrix (zepdrix):

Understand how to solve your problem now or is still confusing? :O

OpenStudy (anonymous):

still confusing haha

OpenStudy (anonymous):

so i plug in and get u(t - 3)f(t - 3)

zepdrix (zepdrix):

\[\large \mathcal L^{-1}\left[e^{-as}\cdot F(s)\right]\quad =\quad \large \mathcal L^{-1}\left[e^{-3s}\cdot\frac{1}{s^2}\right]\] So we just need to think back to what the inverse laplace of 1/s^2 is.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so F(s) = 1/s^2 * e^(-3s)

zepdrix (zepdrix):

No, F(s) is just the 1/s^2 part. See how I tried to match them up in the last equation? e^-3s is our e^-as part of the equation.

zepdrix (zepdrix):

Remember inverse laplace of 1/s^2? :D Look back at your chart if not!! Nice easy one!

OpenStudy (anonymous):

okay so its t

zepdrix (zepdrix):

Yesss good :D So now the exponential term is applying some type of shift. Hmm I wish I had a better understanding of this so I could explain it more thoroughly lol.\[\large t \qquad \rightarrow \qquad (t-3)\;u(t-3)\]

OpenStudy (anonymous):

okay so why isnt it u(t - 3)f(t - 3)

OpenStudy (anonymous):

oh cause its just t

zepdrix (zepdrix):

We were able to find f(t),\[\large f(t)=t\]Our final answer (because of what the exponential is doing to the function) should look like this,\[\large f(t-a)\;u(t-a)\] \[\text{If}\; f(t)=t,\; \text{then}\; f(t-a)=t-a\]

zepdrix (zepdrix):

Hah yah D: grr you too fast lol

OpenStudy (anonymous):

haha okay so then where do i go from there, integration?

zepdrix (zepdrix):

Umm that IS your final answer. Unless your teacher wants it written without the unit step, in which case you might be able to go a step further as write the function as a piece-wise. Hmm.

OpenStudy (anonymous):

oh yeah thats it SWEET

OpenStudy (anonymous):

THANK YOU KINDLY ♥

zepdrix (zepdrix):

Welcome :3